a gas has a volume of 0.785L at 20.1 celsius and 1.00atm. What is the volume in liters of this gas at a temperature of 121 celsius and a pressure of 1.80 atm?
Answer #1

Volume of the gas=0.785L

Temperature =293.25K

Pressure =1.00atm

According to the Gas law that, PV=nRT

Gas Constant = 0.082 Latm mol-1 K-1

So number of moles =1*0.785/293.25*0.082=0.033 moles

At 1.8atm and 394.15K

Volume V=nRT/P = 0.033*0.082*394.15/1.8 = 0.592L